# thanks: Present address

### v4

@@ Abstract
This study investigates the phenomenon of the early-stage impact dynamics in dense suspensions. While Stokes flow theory traditionally applies to suspensions with micrometer-sized particles due to their low Reynolds numbers, suspensions containing larger particles defy such idealizations. Our work focuses on the early-stage impact dynamics in suspensions containing millimeter-sized particles through experiments of impact dynamics. We are particularly interested in the maximum drag force $F_{\mathrm{max}}$ acting on the projectile as a function of the impact speed $u_{0}$. We have successfully conducted experiments using these suspensions and confirmed the relation $F_{\mathrm{max}}\sim u_{0}^{3/2}$ for relatively large $u_{0}$ as observed in the previous studies of suspensions of micrometer-sized particles. Our findings reveal that the early-stage behaviors of millimeter-sized particle suspensions align well with predictions from the floating model, which is equivalent to the viscous model, typically applicable under Stokes flow conditions. This research sheds light on the impact dynamics in dense suspensions, particularly with larger particles, advancing our understanding beyond conventional micrometer-sized systems.
@@ I Introduction
The investigation of impact dynamics within dense suspensions has garnered substantial interest due to its broad applicability across industrial processes and planetary sciences [1]. An intriguing phenomenon within dense suspensions is impact-induced hardening (IIH), enabling locomotion to stop liquids while causing sinking for individuals standing or walking [2]. This behavior finds practical utility in applications such as protective vests [3]. Extensive experimental [4, 5, 6, 7, 8, 9] and theoretical [10, 11, 12] studies have explored IIH, distinguishing it from discontinuous shear thickening (DST) [2, 13, 14, 15] by its localized nature versus DST’s global manifestation, transient behavior versus steadiness, and distinct shear stress responses [10].
@@
A primary focus in impact dynamics research involves elucidating the impact experiments of projectiles[16, 17, 18], where the time evolution of projectiles can be observed to determine the force based on contact mechanics[19]. Early-stage impact dynamics in dense suspensions is governed initially by a viscous force [9, 11, 12], and subsequently, an elastic force plays an important role [8, 10, 11, 12]. Notably, empirical observations suggest a relationship between the maximum drag force $F_{\mathrm{max}}$ and impact speed $u_{0}$, approximated by $F_{\mathrm{max}}\sim u_{0}^{3/2}$ predicted by the viscous model [9] and the floating model [11, 12] which are equivalent models. This result is similar to the relation $F_{\mathrm{max}}\sim u_{0}^{4/3}$ observed for impact processes in dry granular materials [20, 21].
@@
The previous theoretical studies [10, 11, 12] rely on a lattice-Boltzmann method (LBM) and a discrete element method. Although their model analysis agrees with that of experimental results [9], there are various questions about the applicability of their analysis. Their approach is believed to be valid only if the suspended particles are immersed in a viscous fluid in which the Reynolds number is quite low and the Stokes approximation is valid. Because of the limitation of their computer resources, their simulation contains only less than 3000 suspension particles in fluids, and the radius of the projectile is only a few times larger than that of the suspension particles. If suspended particles are a typical suspension, such as cornstarch or potato starch, the projectile must be quite small. Increasing the size of the projectile while maintaining a constant size ratio between the suspension particles and the projectile may invalidate the Stokes approximation at the millimeter scale. This observation highlights the importance of the size scale of the dispersed particles in understanding impact processes at this scale.
@@
Building upon prior investigation, this study experimentally demonstrates how the size scale of dispersed particles influences the early-stage dynamics of the impact processes in dense suspensions. Our exploration delves deeper into the mechanical response of suspension materials to sudden mechanical stimuli, uncovering the fundamental physical mechanisms. We critically reassess the applicability of theoretical analyses utilizing the floating model and LBM within our experimental framework, mainly focusing on early-stage dynamics and the relationship between $F_{\mathrm{max}}$ and $u_{0}$. Our findings affirm that the floating model or the viscous model, typically valid under Stokes flow conditions, holds relevance even for dense suspensions containing millimeter-sized particles.
@@
The rest of the paper is structured as follows. In the next section, we depict the experimental setup. We introduce the theoretical framework, the floating model in detail in Sec. III. We explain the experimental results to clarify the applicability of the floating model in Sec. IV. In Sec. V we discuss our results. We conclude our results with some remarks in Sec. VI. In Appendix A, we clarify the sidewall effect used in the main text from the comparison of the experimental result with a cube container. In Appendix B, we examine whether the simplified buoyancy force can be used to describe experimental results. In Appendix C, we present the exact solution of the floating model.
@@ II Experiment
Let us explain the setup of our experiment. We prepare a dense suspension by mixing millimeter-sized bi-disperse particles, consisting of two types of airsoft pellets: one with a maximum diameter of $d_{\rm max}=8.0$ mm and a density of $1.03~{}{\rm g/cm^{3}}$ (MARUSHIN KOGYO CO.), and the other with a minimum diameter of $d_{\rm min}=6.0$ mm and a density of $1.04~{}{\rm g/cm^{3}}$ (Tokyo Marui Co.), in a NaCl solution whose density is nearly matched with that of the dispersed particles ($\rho_{\rm f}\approx 1.05~{}{\rm g/cm^{3}}$). To observe the impact processes, we primarily use suspensions confined in a quasi-two-dimensional container ( Kenis, Ltd.) with the width $W=123{\rm mm}$, height $H=112{\rm mm}$, and depth $D=14{\rm mm}$.   [^1] As shown in Appendix A, we also examine a cube system with $W=D=94$mm, $H=90$mm to study the boundary effects in the quasi-two-dimensional container. See Fig. 1 as an illustration of the experimental setup.
@@
We introduce the designed packing fraction defined as $\varphi:=V_{\rm p}/V_{\rm s}$ to characterize the density of $N$ number of each particle, where $V_{\rm p}:=\frac{4}{3}\pi N\left[\left(d_{\rm max}/2\right)^{2}+\left(d_{\rm min}/2\right)^{2}\right]$ and $V_{\rm s}:=W\times D\times H$. We prepare the suspension liquid in four different designed packing fractions $\varphi=0.40,0.48,0.51,$ and $0.56$, which correspond to the number of particles $N=243,291,310,$ and $338$ respectively. Note that particles are not uniformly distributed, and most of them float to form a densely packed cluster due to the incomplete density matching. We use the effective packing fraction $\phi$ in the floating cluster near the surface of water for our data analysis in Sec. IV, where $\varphi=0.40,0.48,0.51$, and $0.56$ correspond to $\phi=0.52,0.53,0.54$, and $0.56$, respectively.
@@
We use a metallic sphere as the projectile (diameter $D_{\rm I}=12$ mm, $\rho_{\rm I}=7800~{}{\rm kg\cdot m^{-3}}$), which is suspended by an electromagnet (ESCO Co., Ltd., EA984CM-1) above the suspension before impact. After switching off the magnetic force, the projectile is dropped into the suspension with an initial impact speed $u_{0}$. The impact process is recorded using a high-speed camera (Phantom V641, Phantom T1340, Vision Research) at a frame rate of 10,000 fps. The impact velocity is controlled by varying the release height of the projectile ($h=4\sim 720~{}{\rm mm}$). We conduct repeated impact experiments (10 to 16 times per height) to minimize specific responses from peculiar configurations of the suspended particles.
@@
We extract the trajectory of the projectile by an optical tracking technique with the aid of the Open CV library of Python. We set the marker on the top of the projectile to get its position in each flame. By differentiating the time-evolution of the position with respect to time, we evaluate the time-evolutions of the velocity and the acceleration of the projectile. Subtracting the buoyancy force acting on the projectile from the equation of motion of the projectile, we can evaluate the drag force acting on the projectile.
@@ III Floating model
Brassard et al. [9] proposed the viscous model to describe the motion of the projectile in dense suspensions of micron-meter-sized particles in the early stage of the impact dynamics. Later, Pradipto and Hayakawa [11, 12] renamed it as the floating model, adding its derivation. To get IIH, we need an elastic force [6, 7, 8, 10, 11, 12], which appears only in a relatively late stage of the impact process, although the floating model involves only the viscous drag force acting on the projectile, except for the buoyancy force. In this paper, we adopt the floating model to describe the motion around $t_{\mathrm{max}}$ at which the drag force takes $F_{\mathrm{max}}$. We depict an impact process in a suspension in the left figure of Fig. 2, where the position of the bottom head of the projectile with radius $a_{\rm I}$ and its density $\rho_{\rm I}$ is denoted by $z$ and $z=0$ is fixed on the surface of the suspension. Thus, we are only interested in the position $z$ satisfying $z\leq 0$, i,e, under the influence of the suspension liquid. As a result of the drag force $F_{\rm D}$ and buoyancy force $F_{\mathrm{B}}$ acting on the projectile, the projectile speed decreases with time and approaches zero. We should note that the maximum force $F_{\mathrm{max}}$ acting on the projectile appears in the relatively early stage of impact processes, and the behavior of the projectile around $t_{\mathrm{max}}$ can be understood without the elastic force acting on the projectile, at least, for micro-meter sized suspensions and LBM simulations [9, 11, 12].
Figure 2: An illustration of a projectile in a suspension liquid, where the impact speed is $u_{0}$ when the projectile attaches to the surface of the liquid (Left). We denote the position of the bottom head of the projectile $z(<0)$ where the center of mass is located at $z_{\rm I}$, the radius of the projectile $a_{\rm I}$, and the density of it is $\rho_{\rm I}$. The density of the solvent is $\rho_{\rm f}$. We introduce the effective viscosity $\eta_{\mathrm{eff}}$ acting on the projectile as a mean-field viscosity of the suspension liquid in the floating model (Right).
@@
$F_{\rm B}\simeq-m_{\rm I}g\left(1-\frac{\rho_{\rm f}}{\rho_{\rm I}}\right).$ (1)
where $m_{\rm I}$ is the mass of the projectile, $g$ is the gravitational constant, $\rho_{\rm f}$ is the density of the solvent and $\rho_{\rm I}$ is the density of the projectile. The validity of this simplification is argued in Appendix B.
@@
$F_{\rm D}=3\pi\eta_{{\rm eff}}\frac{dz}{dt}z$ (2)
where $\eta_{\rm eff}$ is an effective viscosity. For simplicity, we assume that $\eta_{\mathrm{eff}}$ is independent of $u_{0}$ but depends on $\phi$. The expression of Eq. (2) is reduced to the Stokes drag force for a filled projectile in the liquid for $z\leq-2a_{\rm I}$. Thus, Eq. (2) is only valid for $-2a_{\rm I}<z<0$. It is controversial whether the drag force $F_{\mathrm{D}}$ can be used for a suspension of millimeter-sized particles because fluid flows around a large particle to generate vortices. Nevertheless, we may examine Eq. (2) for a fluid flow around a projectile in a dense suspension of millimeter-sized particles because the interstitial distance between suspended particles is too small to generate vortices. The validity of this model will be examined by the comparison between the model analysis and experimental results.
@@
$\frac{d^{2}Z}{d\tau^{2}}=-\left(1-\xi\right)+\eta\frac{dZ}{d\tau}Z$ (4)
@@
$\tilde{F}_{\mathrm{max}}=\frac{\sqrt{2\eta}}{3}U_{0}^{\frac{3}{2}},\qquad\left(\eta U_{0}\gg 1\right)$ (6)
@@ IV Results
Figure 3: A set of experimental snapshots of an impact process for $\phi=0.56$ and $u_{0}=3.1~{}{\rm m/s}$. Note that 2.1 ms after the first impact (the middle figure) the force acting on the projectile exhibits the maximum value $F_{\mathrm{max}}$. The movie is available at Ref.[22].
@@
Let us present the experimental results and examine whether the floating model introduced in the previous section can be used in the early stage of dynamics focusing on the relationship between $F_{\mathrm{max}}$ and $u_{0}$. Figure 3 is a typical set of experimental snapshots of an impact process for $\varphi=\phi=0.56$ and $u_{0}=3.1~{}{\rm m/s}$. The impact speed $u_{0}$ and the initiation time of the impact ($t=0$) are defined as the points of its maximum velocity and the corresponding time, respectively. Then, the time $t$ begins with the initiation time. Figure 3 (b) is a snapshot corresponding to the first hit of the projectile on the liquid surface at $t=0$. Figure  3 (c) is a snapshot when the drag force acting on the projectile reaches the maximum value $F_{\mathrm{max}}$ at $t_{\mathrm{max}}=2.1$ ms after the first impact. The movie of the impact process is available in Ref. [22].
@@
Figure 4: The time evolution of the velocity of the projectile (a) and the drag force acting on the projectile (b) for $\phi=0.56$ and $u_{0}=3.1~{}{\rm m/s}$. The blue dots are experimental data. The solid lines are the solution of Eq. (3) with a fitting parameter $\eta$. The dashed vertical red line indicates the time $t_{\mathrm{max}}$ to take $F_{\mathrm{max}}$.
@@
The projectile stops in the middle of the suspension for the dense suspensions ($\phi\geq 0.54$ corresponding to $\varphi\geq 0.51$), where the projectile cannot sneak in the suspension liquid for low-speed impact and it can sneak in the liquid completely for high-speed impact. For lower packing fractions such as $\phi=0.52$ (corresponding to $\varphi=0.40$), the projectile penetrates the liquid completely to reach the bottom of the container.
@@
Figure 5: The plots between the dimensionless maximum force $\tilde{F}_{\mathrm{max}}:=F_{\mathrm{max}}/m_{\rm I}g$ and impact-speed $U_{0}:=u_{0}/\sqrt{a_{\rm I}g}$ for various packing fractions $\phi=0.52$ (a), 0.53 (b), 0.54 (c), and 0.56 (d) corresponding to $\varphi=0.40,0.48,0.51$, and $0.56$, respectively. The blue solid circles indicate the experimental results and the black solid lines are obtained by the numerical solutions of the floating model Eq. (4). The gray area around the solid line indicates the possible solutions within the errors of estimated $\eta$. The black solid lines in large $U_{0}$ are the guide lines $\tilde{F}_{\mathrm{max}}\propto U_{0}^{3/2}$ as expected from Eq. (6).
@@
As explained in Sec. II we can trace the velocity and acceleration of the projectile. By using Eq. (1) we can evaluate the drag force acting on the projectile. Figures 4 depict the time evolutions of the velocity and the drag force acting on the projectile. As shown in Fig. 4 the projectile slows down after the impact by the drag force $F_{D}$ where $F_{D}$ takes the maximum value $F_{\mathrm{max}}$ at $t_{\mathrm{max}}$. In these figures, the dots are obtained in the experiments, and the solid lines are the fitting curves using the solution of Eq.(3). We choose $\eta$ as a fitting parameter to recover experimental $\tilde{F}_{\mathrm{max}}$ in the floating The fitting is performed by obtaining $\eta$ to give the observed in the experiment. Surprisingly, the solution of the floating model Eq. (3) with a fitting parameter $\eta:=3\pi\eta_{{\rm eff}}a_{\rm I}^{3/2}/(g^{1/2}m_{\rm I})$ gives reasonable agreement with the experimental results, at least, for $t<t_{\mathrm{max}}\approx 2.1$ ms corresponding to $F_{\mathrm{max}}$, although there is some discrepancy between the solution of the floating model and the experimental results for $t>t_{\mathrm{max}}$, in which an elastic force might be involved. Here we focus only on the early stage dynamics for $t\leq t_{\mathrm{max}}$ in this paper, the late stage falls outside the scope of the present study. Thus, this result suggests that the floating model is valid to describe the motion of the projectile in the early stage, for $t\leq t_{\mathrm{max}}$, $u_{0}=3.1$ m/s, and $\phi=0.56$.

### v5

# Early-stage impact dynamics in dense suspensions of millimeter-sized particles

@@ Abstract
We investigate the early-stage impact dynamics in dense suspensions composed of millimeter-sized particles. While traditional models based on Stokes flow are typically applicable to suspensions of micrometer-sized particles, their validity for larger particles remains uncertain. Through controlled impact experiments, we examine the maximum drag force $F_{\mathrm{max}}$ acting on a projectile as a function of impact speed $u_{0}$. We have successfully conducted experiments using these suspensions and confirmed the relation $F_{\mathrm{max}}\sim u_{0}^{3/2}$ for relatively large $u_{0}$ as observed in the previous studies of suspensions of micrometer-sized particles. Our results demonstrate that the floating model—equivalent to the viscous model—successfully captures the early-stage dynamics even in suspensions with larger particles. This finding suggests that viscous-dominated behavior persists under certain conditions, extending the applicability of the floating model or the viscous model to new regimes. Our work provides experimental validation for theoretical predictions and contributes to a deeper understanding of impact-induced responses in dense suspensions.
@@ I Introduction
Suspensions of colloidal particles in liquid can be commonly observed in nature [1, 2, 3]. Knowing the flow properties of a suspension is important for both natural science and industry. Some suspensions, as typical non-Newtonian fluids, exhibit discontinuous shear thickening (DST) in which the viscosity abruptly changes discontinuously from a small value to a large value at a critical shear rate. DST has attracted considerable attention among physicists in recent years  [4, 5, 3, 6, 7, 8, 9, 10] as a typical nonequilibrium discontinuous phase transition between a liquid-like phase and a solid-like phase. The DST is also expected to be important for industrial applications such as traction controls. There are various spectacular aspects in addition to DST in the rheology of dense suspensions. See, e.g., Refs. [11, 12, 13, 14, 15] for both experimental and theoretical aspects.
An intriguing phenomenon within dense suspensions is impact-induced hardening (IIH), enabling locomotion to stop liquids while causing sinking for individuals standing or walking [8]. This is related to the impact dynamics of a projectile on a plate [16, 17, 18, 19, 20, 21], or a solid particle [16, 22, 23, 24, 25, 26], or granular beds [27, 28, 29, 30, 31, 32, 33], but the impact dynamics in suspensions is more drastic because IIH is related to a dynamical phase transition from a liquid to a solid. This behavior finds practical utility in applications such as protective vests [34]. Extensive experimental [35, 36, 37, 38, 39, 40, 41] and theoretical [42, 43, 44] studies have explored IIH including oscillation and stick-slip processes [45] and relaxation process after the impact [46, 47, 48]. They found that IIH is caused by an elastic contact force between suspended particles, mostly observed only in suspensions of frictional particles confined in a shallow container [40, 43]. Some papers confirmed the existence of the dynamical jammed region in front of the projectile [35, 36, 37, 44]. They also confirmed that there is a viscous regime in an early stage of the impact process before the reaching the maximum drag force acting on the projectile $F_{\mathrm{max}}$ at time $t_{\mathrm{max}}$ [41, 43, 44], and subsequently, an elastic force plays an important role [40, 42, 43, 44]. Although IIH is similar to DST, IIH can be distinguished from DST [8, 49, 6, 10] by its localized nature versus the global manifestation of DST, transient behavior versus steadiness, and distinct shear stress responses [42].
@@
A primary focus in impact dynamics research involves elucidating the impact experiments of projectiles[17, 19, 20, 21], where the time evolution of projectiles can be observed to determine the force based on contact mechanics[50]. Notably, empirical observations suggest a relationship between the maximum drag force $F_{\mathrm{max}}$ and impact speed $u_{0}$, approximated by $F_{\mathrm{max}}\sim u_{0}^{3/2}$ predicted by the viscous model [41] and the floating model [43, 44], which are equivalent models. This result is similar to the relation $F_{\mathrm{max}}\sim u_{0}^{4/3}$ observed for impact processes in dry granular materials [33, 51].
@@
The previous numerical studies [42, 43, 44] rely on a lattice-Boltzmann method (LBM) and a discrete element method. Although their model analysis agrees with the experimental results [41], several questions remain about the applicability of their analysis. Their approach is believed to be valid only if the suspended particles immersed in a viscous fluid are sufficiently small that the Reynolds number is quite low and the Stokes approximation is valid. Due to the limitations of their computer resources, their simulation contains fewer than 3000 suspension particles in fluids, and the radius of the projectile is only a few times larger than that of the suspension particles. If suspended particles are a typical suspension, such as cornstarch or potato starch, the projectile must be quite small. Increasing the size of the projectile while maintaining a constant size ratio between the suspension particles and the projectile may invalidate the Stokes approximation at the millimeter scale. This observation highlights the importance of the size of the dispersed particles in understanding impact processes at this scale.
@@
Building upon prior investigation, this study experimentally demonstrates how the size of dispersed particles influences the early-stage dynamics of the impact processes in dense suspensions. Our exploration delves deeper into the mechanical response of suspension materials to sudden mechanical stimuli, uncovering the fundamental physical mechanisms. We critically reassess the applicability of theoretical analyses utilizing the floating model and LBM within our experimental framework, mainly focusing on early-stage dynamics and the relationship between $F_{\mathrm{max}}$ and $u_{0}$. Our findings affirm that the floating model or the viscous model, typically valid under Stokes flow conditions, holds relevance even for dense suspensions containing millimeter-sized particles. The dramatic IIH process in the impact dynamics caused by an elastic response of suspensions is not the target of this study, but will be the subject in the proceeding research.
@@
The rest of the paper is structured as follows. In the next section, we depict the experimental setup. We introduce the theoretical framework, the floating model in detail in Sec. III. We explain the experimental results to clarify the applicability of the floating model in Sec. IV. In Sec. V we discuss our results. We conclude our results with some remarks in Sec. VI. In Appendix A, we clarify the sidewall effect used in the main text from the comparison of the experimental result with a cube container. In Appendix B, we examine whether the simplified buoyancy force can be used to describe experimental results. In Appendix C, we present the exact solution of the floating model. In Appendix D, we examine how the experimental results depend on the projectile diameter and the viscosity of the solvent. In Appendix E, we present a phenomenology to derive the effective viscosity for dense suspensions.
@@ II Experiment
Let us explain the setup of our experiment. We prepare a dense suspension by mixing millimeter-sized bi-disperse particles, consisting of two types of airsoft pellets: one with a maximum diameter of $d_{\rm max}=8.0$ mm and a density of $1.03~{\rm g/cm^{3}}$ (MARUSHIN KOGYO CO.), and the other with a minimum diameter of $d_{\rm min}=6.0$ mm and a density of $1.04~{\rm g/cm^{3}}$ (Tokyo Marui Co.), in a NaCl solution in the main text whose density is nearly matched with that of the dispersed particles ($\rho_{\rm f}\approx 1.05~{\rm g/cm^{3}}$). To observe the impact processes, we primarily use suspensions confined in a quasi-two-dimensional container ( Kenis, Ltd.) with the width $W=123{\rm mm}$, height $H=112{\rm mm}$, and depth $D=14{\rm mm}$. See Fig. 1 as an illustration of the experimental setup. As shown in Appendix A, we also examine a cube system with W=D=94W=D=94mm, H=90H=90mm to clarify the boundary effects in the quasi-two-dimensional container. Then, we have confirmed that the results depend little on the side boundaries or box shapes.
@@
We introduce the designed packing fraction defined as $\varphi:=V_{\rm p}/V_{\rm s}$ to characterize the density of $N$ number of each particle, where $V_{\rm p}:=\frac{4}{3}\pi N\left[\left(d_{\rm max}/2\right)^{2}+\left(d_{\rm min}/2\right)^{2}\right]$ and $V_{\rm s}:=W\times D\times H$. We prepare the suspension liquid in four different designed packing fractions $\varphi=0.40,0.48,0.51,$ and $0.56$, which correspond to the number of particles $N=243,291,310,$ and $338$ respectively. Note that $\varphi=0.56$ is the maximum possible packing fraction as the suspension is no longer slurry [52] if we study denser situations. Note that particles are not uniformly distributed, and most of them tend to form a densely packed cluster due to imperfect density matching. We use the effective packing fraction $\Phi$ in the floating cluster near the surface of water for our data analysis in Sec. IV, where $\varphi=0.40,0.48,0.51$, and $0.56$ correspond to $\Phi=0.52,0.53,0.54$, and $0.56$, respectively. Although the dynamical jammed region produced by an impact becomes denser than the initial $\Phi$, the amount of compactification is not large [44]. Also, the experimental determination of the packing fraction in the dynamical jammed region is difficult. Therefore, we use only $\Phi$ for the analysis of experimental results.
@@
We use a metallic sphere as the projectile (diameter $D_{\rm I}=12$ $m$, $\rho_{\rm I}=7800~{\rm kg\cdot m^{-3}}$ in the main text), which is suspended by an electromagnet (ESCO Co., Ltd., EA984CM-1) above the suspension before impact. After switching off the magnetic force, the projectile is dropped into the suspension with an initial impact speed $u_{0}$. The impact process is recorded using a high-speed camera (Phantom V641, Phantom T1340, Vision Research) at a frame rate of 10,000 fps. The impact velocity is controlled by varying the release height of the projectile ($h=4\sim 720~{\rm mm}$). We conduct repeated impact experiments (10 to 16 times per height) to minimize specific responses from peculiar configurations of the suspended particles.
@@
We extract the trajectory of the projectile by an optical tracking technique with the aid of the Open CV library of Python. We set the marker on the top of the projectile to get its position in each flame. We adopt the second-order central difference scheme to obtain the velocity and the acceleration of the projectile. Subtracting the buoyancy force acting on the projectile from the equation of motion of the projectile, we can evaluate the drag force acting on the projectile.
@@ III Floating model
Brassard et al. [41] proposed the viscous model to describe the motion of the projectile in dense suspensions of micron-sized particles in the early stage of the impact dynamics. Later, Pradipto and Hayakawa [43, 44] renamed it as the floating model, adding its derivation. [^1] This name originates from floating suspended particles that do not have any elastic response even if a projectile hits suspended particles because they are floating. To get IIH, we need an elastic force [38, 39, 40, 42, 43, 44], which appears only in a relatively late stage of the impact process, although the floating model involves only the viscous drag force acting on the projectile, except for the buoyancy force. In this paper, we adopt the floating model to describe the motion around $t_{\mathrm{max}}$ at which the drag force takes $F_{\mathrm{max}}$. We depict an impact process in a suspension in the left figure of Fig. 2, where the position of the bottom head of the projectile with radius $a_{\rm I}:=D_{\mathrm{I}}/2$ and its density $\rho_{\rm I}$ is denoted by $z$ and $z=0$ is fixed on the surface of the suspension. Thus, we are only interested in the position $z$ satisfying $z\leq 0$, i,e, under the influence of the suspension liquid. As a result of the drag force $F_{\rm D}$ and buoyancy force $F_{\mathrm{B}}$ acting on the projectile, the projectile speed decreases with time and approaches zero. We should note that the maximum force $F_{\mathrm{max}}$ acting on the projectile appears in the relatively early stage of impact processes, and the behavior of the projectile around $t_{\mathrm{max}}$ can be understood without the elastic force acting on the projectile, at least, for micro-meter sized suspensions and LBM simulations [41, 43, 44].
Figure 2: An illustration of a projectile in a suspension liquid, where the impact speed is $u_{0}$ when the projectile attaches to the surface of the liquid (Left). We denote the position of the bottom head of the projectile $z(<0)$ where the center of mass is located at $z_{\rm I}$, the radius of the projectile $a_{\rm I}$, and the density of it is $\rho_{\rm I}$. The density of the solvent is $\rho_{\rm f}$. We introduce the effective viscosity $\eta_{\mathrm{eff}}$ acting on the projectile as a mean-field fluid of the suspension liquid in the floating model (Right).
@@
$F_{\rm B}\approx-m_{\rm I}g\left(1-\frac{\rho_{\rm f}}{\rho_{\rm I}}\right),$ (1)
where $m_{\rm I}$ is the mass of the projectile, $g$ is the gravitational acceleration, $\rho_{\rm f}$ is the density of the solvent and $\rho_{\rm I}$ is the density of the projectile. The validity of this simplification is argued in Appendix B.
@@
$F_{\rm D}=3\pi\eta_{{\rm eff}}\frac{dz}{dt}z,$ (2)
where $z$ is the vertical position of the head (deepest position) of the projectile, and $\eta_{\rm eff}$ is an effective viscosity of the mean-field fluid (see Fig. 2). For simplicity, we assume that $\eta_{\mathrm{eff}}$ is independent of $u_{0}$ but depends on $\Phi$. Although $\eta_{\mathrm{eff}}$ is the mean-field viscosity of the dynamical jammed region [44] corresponding to the growth of the dynamical jammed region [36], it will be treated as a fitting parameter in the later analysis because of the difficulty of estimating the local viscosity. The expression of Eq. (2) is reduced to the Stokes drag force for a filled projectile in the liquid for $z\leq-2a_{\rm I}$. Thus, Eq. (2) is only valid for $2a_{\rm I}<z<0$.
The validity of the drag force in Eq. (2) has already been confirmed for the impact dynamics in standard suspensions, such as cornstarch and potato-starch, and microscopic simulations [41, 43, 44]. It is, however, controversial whether the drag force $F_{\mathrm{D}}$ can be used for a suspension of millimeter-sized particles because fluid flows around a large particle to generate vortices. Nevertheless, we may examine Eq. (2) for a fluid flow around a projectile in a dense suspension of millimeter-sized particles because the interstitial distance between suspended particles is too small to generate vortices. The validity of this model will be examined by comparing the model analysis and experimental results.
@@
$\frac{d^{2}Z}{d\tau^{2}}=-\left(1-\xi\right)+\eta\frac{dZ}{d\tau}Z,$ (4)
@@
$\tilde{F}_{\mathrm{max}}\approx\frac{\sqrt{2\eta}}{3}U_{0}^{\frac{3}{2}},\qquad\left(\eta U_{0}\gg 1\right)$ (6)
@@ IV Results
Figure 3: A set of experimental snapshots of an impact process for $\Phi=0.56$ and $u_{0}=1.6~{\rm m/s}$. Note that 3.0 ms after the first impact (the middle figure) the force acting on the projectile exhibits the maximum value $F_{\mathrm{max}}$ (Multimedia available online).
@@
Let us present the experimental results and examine whether the floating model introduced in the previous section can be used in the early stage of dynamics, focusing on the relationship between $F_{\mathrm{max}}$ and $u_{0}$. Figure 3 is a typical set of experimental snapshots of an impact process for $\varphi=\Phi=0.56$ and $u_{0}=1.6~{\rm m/s}$ (Multimedia available online). The impact speed $u_{0}$ and the initiation time of the impact ($t=0$) are defined as the points of its maximum velocity and the corresponding time, respectively. Then, the time $t$ begins with the initiation time. Figure 3 (b) is a snapshot corresponding to the first hit of the projectile on the liquid surface at $t=0$. Figure  3 (c) is a snapshot when the drag force acting on the projectile reaches the maximum value $F_{\mathrm{max}}$ at $t_{\mathrm{max}}=3.0$ ms after the first impact.
Interestingly, most trials exhibit that the projectile hits one of the suspended particles at the first hit on the liquid surface. This might cause some elastic response. However, since the suspended particles are floating and particles are lubricated with the other particles, the hit of the projectile suppresses elastic response (Fig. 4). If the suspended particles are percolated from the contacting particles with the projectile to the bottom plate, we can get a significant elastic response [40, 43]. However, this only appears in the late stage for $t>t_{\mathrm{max}}$. Nevertheless, there are elastic effects caused by contact between the projectile and suspended particles that are not accounted for in the floating model.
Figure 4: The floating relaxation of dispersed particles after an impact, where a black ball represents the projectile. After the impact, the floating suspended particles are moved to relax into another stable position.
@@
Figure 5: The time evolution of the velocity of the projectile (a) and the drag force acting on the projectile (b) for $\Phi=0.56$ and $u_{0}=1.6~{\rm m/s}$. The blue dots are experimental data. The solid lines are the solution of Eq. (3) with a fitting parameter $\eta$. The dashed vertical red line indicates the time $t_{\mathrm{max}}$ to take $F_{\mathrm{max}}$.
@@
The projectile stops in the middle of the suspension for the dense suspensions ($\Phi\geq 0.54$ corresponding to $\varphi\geq 0.51$), where the projectile cannot sneak into the suspension liquid for low-speed impact, and it can sneak into the liquid completely for high-speed impact. For lower packing fractions, such as $\Phi=0.52$ (corresponding to $\varphi=0.40$), the projectile penetrates the liquid completely and reaches the bottom of the container.
@@
Figure 6: The plots between the dimensionless maximum force $\tilde{F}_{\mathrm{max}}:=F_{\mathrm{max}}/m_{\rm I}g$ and impact-speed $U_{0}:=u_{0}/\sqrt{a_{\rm I}g}$ for various packing fractions $\Phi=0.52$ (a), 0.53 (b), 0.54 (c), and 0.56 (d) corresponding to $\varphi=0.40,0.48,0.51$, and $0.56$, respectively. The blue solid circles indicate the experimental results, and the dark-gray solid lines are obtained by the numerical solutions of the floating model Eq. (4). The gray area around the solid line indicates the possible solutions within the errors of the estimated $\eta$. The black solid line segments in large $U_{0}$ are the guide lines $\tilde{F}_{\mathrm{max}}\propto U_{0}^{3/2}$ as expected from Eq. (6).
@@
As explained in Sec. II we can trace the velocity and acceleration of the projectile. By using Eq. (1) we can evaluate the drag force acting on the projectile. Figures 5 depict the time evolutions of the velocity and the drag force acting on the projectile. As shown in Fig. 5 the projectile slows down after the impact by the drag force $F_{D}$ where $F_{D}$ takes the maximum value $F_{\mathrm{max}}$ at $t_{\mathrm{max}}$. In these figures, the dots are obtained in the experiments, and the solid lines are the fitting curves using the solution of Eq.(3). We choose $\eta:=3\pi\eta_{{\rm eff}}a_{\rm I}^{3/2}/(g^{1/2}m_{\rm I})$ as a fitting parameter to recover experimental $\tilde{F}_{\mathrm{max}}$ in the floating model. The behavior of the experimental data is qualitatively similar to that obtained from the floating model, although the very early stage dynamics of the projectile observed in the experiment seem to differ from the prediction of the floating model. Therefore, we cannot expect quantitative agreement between theoretical $t_{\mathrm{max}}$ and experimental $t_{\mathrm{max}}$ as shown in Fig. 5. This disagreement between the theory and the experiment differs from the agreement between the LBM simulation and the theory [43, 44]. We also stress that the experimental data exhibit different time evolution from that of the floating model for $t>t_{\mathrm{max}}$. Here, we focus only on the early stage dynamics for $t\leq t_{\mathrm{max}}$ in this paper; the late stage falls outside the scope of the present study.